OpenStudy (anonymous):

The eighth grade class received surveys regarding internet access in their homes. Three-fourths of the eighth grade class has access to the internet in their homes. If 120 students have access to the internet, how many students are in the eighth grade?

OpenStudy (anonymous):

no of students are in the eighth grade= 4/3 *120 =160

OpenStudy (ghazi):

lets say there are x students then x/3 has internet access which is equal to 120 x/3=120 x=120*3

OpenStudy (ghazi):

sorry its 3/4x=120

OpenStudy (ghazi):

\[\frac{ 3 }{ 4 }x=120\] x=120*4/3= 480/3=160

OpenStudy (anonymous):

Thank you, that help was much needed.

OpenStudy (ghazi):

:)